OpenStudy (anonymous):

the equation a= 180(n-2)/n represents the angle measures, a, in a regular n-sided polygon. When the equation is solved for n, n is equal to a fraction with a denominator of a – 180. What is the numerator of the fraction?

OpenStudy (welshfella):

a = angle measures Do you mean the measure of one angle in the polygon or the total measure of all the angles?

OpenStudy (anonymous):

the equation a= 180(n-2)/n represents the angle measures, a, in a regular n-sided polygon. When the equation is solved for n, n is equal to a fraction with a denominator of a – 180. What is the numerator of the fraction?

OpenStudy (anonymous):

sorry i forgot to write the equation in

OpenStudy (welshfella):

an = 180n - 360 an - 180n = -360 can you continue?

OpenStudy (anonymous):

ummm, maybe?

OpenStudy (welshfella):

factor the LHS . Take out n.

OpenStudy (anonymous):

What is LHS?

OpenStudy (welshfella):

the left side of the equation

OpenStudy (anonymous):

Okay, sorry i'm not the smartest... at all

OpenStudy (welshfella):

n is common to both terms

OpenStudy (anonymous):

alright, so do i divide n from both sides?

OpenStudy (welshfella):

no factor an - 180n = n(a-180) = -360 now you divide both sides by a - 180

OpenStudy (welshfella):

you have the numerator on RHS already

OpenStudy (anonymous):

Alright..

OpenStudy (welshfella):